# Incorporating Gradients to Rules: Towards Lightweight, Adaptive Provenance-based Intrusion Detection

### v1

@@ VIII Conclusion
This paper introduces Captain, a tag-based PIDS capable of automatically adapting to diverse detection environments. Captain’s flexibility, enhanced by three adaptive parameters: tag initialization ($A$), propagation rate ($G$), and threshold ($T$), distinguishes it from existing rule-based PIDS. We have developed a differentiable tag propagation framework that records gradients, enabling the optimization of these parameters using the gradient descent algorithm. To our knowledge, this is the first effort to incorporate gradient descent methods in optimizing rule-based PIDS. We then evaluate Captain using several APT attack datasets. Our results demonstrate a remarkable capability of Captain to adapt to different environments automatically, significantly reducing false alarms, detection latency, and runtime overhead, outperforming the SOTA baselines.
@@ Appendix A Attack Scenarios in Datasets
A-A DARPA Engagement
TABLE VIII: Dataset Details
Dataset OS Duration(hh:mm) # of Nodes # of Events Attacks
E3-Trace Linux 263:05 14.10M 20.86M Firefox Backdoor, Pine Backdoor, Phishing Email
E3-Theia Linux 247:02 1.587M 27.47M Firefox Backdoor, Browser Extension, Phishing E-mail
E3-Cadets FreeBSD 263:28 614.5K 8.127M Nginx Backdoor
E5-Trace Linux 248:23 173.6M 1596M Firefox Elevate Inject, Nmap Scan
E5-Cadets FreeBSD 249:05 14.13M 1194M Nginx Backdoor, Nmap Scan
SOC-Streaming Linux 06:51 302.9K 1.731M ReverseShell, WebShell, AttackChain, Kimsuky, Chaos
SOC-Dev Linux 06:32 78.89K 187.1K ReverseShell, WebShell, AttackChain, Kimsuky, Chaos
SOC-Cloud Linux 06:42 2.464M 5.817M ReverseShell, AttackChain, Kimsuky, Chaos
TABLE IX: Training and testing time for each dataset. The dataset durations are in the format ($hh:mm$). The dataset size is denoted by the number of events.
Dataset Training Testing
Dur. Size Dur. Size
T-3 174:55 15.7M 88:10 5.2M
C-3 77:53 1.9M 185:36 5.4M
T-5 70:45 119.6M 177:38 312.0M
C-5 71:15 59.8M 177:50 208.8M
S-1 5:23 8.5M 1:28 3.6M
S-2 5:01 1.2M 1:31 629K
S-3 5:13 209.0K 1:29 113K
E3-Tracecontains five attack attempts. In the first attempt, Firefox on the target host is exploited by visiting a malicious website. The attackers then uploaded a payload and obtained a remote shell with root privilege. We will refer to this attack later as Firefox Backdoor. In the second attempt, the attackers sent a phishing e-mail with a link to a website hosted by the attackers. The victim entered the requested information on the website, and the attackers received the victim’s personal information. This attack only involves network activities, and the malicious part is not visible in the audit data. We only know the victim visited an external website, which is a common activity. Thus, we removed this attack from our evaluation. In the third attempt, the attackers tried to exploit the target with a malicious browser extension but failed to connect to the target. Thus, we didn’t include this attempt in our evaluation. The fourth attempt continues the previous attempt of the malicious browser extension exploitation. This time, the attacker successfully obtained a remote shell. Then, they wrote and executed a malicious executable to perform portscan. We will refer to this attack later as Pine Backdoor. The fifth attempt involves a phishing e-mail with executable attachments. The victim downloaded and executed the malicious executable; hence, the attackers obtained a remote console and performed a port scan on the target host. We will refer to this attack later as Executable Attachment.
Besides Trace, we also analyzed attack campaigns in E3-Cadetsdataset. There are four attack attempts made using the same technique. Three out of four attempts failed, so we only included the successful one in our evaluation. In this attempt, the Nginx web server was exploited with a malformed HTTP request, and the attackers gained access to a remote shell. Then, the attackers downloaded a malicious payload, which connected back to the attackers’ listener to establish a persistent connection. The attackers also used the payload to perform a port scan. We will refer to this attack later as Nginx Backdoor.
A-B Simulated Environments
TABLE X: Simulated Scenario Summary
Attack Scenario Attack Description
Malicious ReverseShell(RS) Connect to the victim’s host, collect system information, and install multiple applications
@@ A-B Simulated Environments
Server Environment App or Process Involved
Benign Streaming Kafka, Mysql, Nginx, Redis, Zookeeper
@@
The simulated APT attack chains are generated with our automated attack simulator. The automated attack generator is derived from the Atomic Red Team project [6], a library of atomic attacks mapped to the MITRE ATT&CK framework. We modified it so we can either determine or randomly select an atomic attack to execute at each attack stage and thus generate an attack chain from the individual tests. When we design the screenplay for simulated attacks, we leave most of the stages to be randomly selected to avoid manually introduced bias.
We designed three kinds of simulated attacks, including fully randomized, partially randomized, and deterministic attacks, to demonstrate the adaptability of our model. Each step in a fully randomized attack is randomly selected from the Atomic Red Team atomic attack library. A partially randomized attack has determined steps at some stages but randomly chooses steps for the rest of the steps. A deterministic attack has defined steps at every attack stage. We used one deterministic attack, the North Korean APT Kimsuky (Kim), and 3 partially randomized attacks, comprising a reverse shell attack (RS), a web shell attack (WS), and a payload attack (Chaos) in our evaluation. Kimsuky is an active APT group based in North Korea. We refer to a threat intelligence report[10] to select techniques used in each attack stage and reproduce North Korean APT Kimsuky with the help of Atomic Red Team[6]. The reverse shell, the web shell, and the payload attack uploaded three different payloads to establish the connection between the target host and the attackers to complete the attack. Finally, we also generated fully randomized attack chains (AC) using Atomic Red Team for the evaluation.
As for the benign scenarios, the SOC provided a list of detailed operation environments in the real world. Since some operation environments have similar configurations, we selected a subset of rather distinguished environments to reproduce. When we select scenarios to simulate, we also intentionally include applications and services that often trigger false alarms. Eventually, we reproduced three representative environments. Streaming simulates a data store and streaming server. Dev simulates a development server. Cloud simulates a cloud server. Table X specifies the apps or processes running in each simulated environment.
Once we have the simulated attacks and operation environments, we conduct the experiments following the attack campaign pattern used in DARPA Engagements to make it as realistic as possible. We start the data collector in three simulated environments and keep it running for a day. During the day, we ask a red team to perform attacks on all three targets at some time. In the end, we have the data produced in a day from 3 hosts and the ground truth of attacks performed. Table VIII summarizes the information of the three datasets we collected in the described manner.
@@ Appendix B Captain Policies
B-A Tag Propagation Policies
We adopted tag propagation policies from Morse [19] in our implementation. The tag propagation policies are detailed in Table XI. In the table, $dtag$ refers to the data tag and $ptag$ refers to the code tag. $T_{qb}$ and $T_{qe}$ stand for quiescent tag values for benign and suspect environment processes. $d_{b}$ and $d_{e}$ are decay rates for benign and suspect environment processes. $a_{b}$ and $a_{e}$ are the attenuation factors for benign and suspect environment subjects. Please refer to the original paper for formal and detailed definitions of these parameters. To get a fair comparison, we set all hyperparameters in Table XI as the default values according to the recommendation in the original paper.
TABLE XI: Tag propagation policies
@@ B-A Tag Propagation Policies
suspect suspect environment
@@
B-B Alarm Generation Policies
We adopted alarm generation policies from Morse [19] in our implementation. The alarm generation policies are detailed in Table XII. $incl\_exec(p)$ means $p$ is contains the execution permission. $socket(o)$ holds when $o$ refers to a socket.
TABLE XII: Alarm generation policies
@@ B-B Alarm Generation Policies
Appendix C Additional Experiment Results
C-A False Alarm Entity Reduction
Notably, in simple scenarios such as S-2 and S-3, the size of the malicious sub-graph is marginally reduced. This is primarily because most malicious activities and related entities in these simulated environments are not seen during the training stage. However, Captain can still reduce the false alarms by adjusting the alarm threshold.
In addition to minimizing false alarms, we also want to reduce the number of suspicious nodes in the provenance graph, thereby expediting forensic investigations. In forensic analysis using tag-based PIDS, it’s essential to examine the subgraph comprising all nodes tagged as malicious to comprehend the attack. Consequently, if malicious tags are widespread throughout the graph, the resulting subgraph may become too large for effective processing and analysis.
Figure 7: The comparison between Captain and the baseline in terms of the numbers of low-integrity nodes in the provenance graph
C-B Amount of Gradients to Save During Training
We maintain two dictionaries for each node $n$ to save the gradients of $tag_{n}$:
$\nabla_{n}^{A}:=\{n^{\prime}:\frac{\partial tag_{n}}{\partial a_{n}^{\prime}},n^{\prime}\in N\}$ and $\nabla_{n}^{G}:=\{e:\frac{\partial tag_{n}}{\partial g_{e}},e\in E\}$
For each node, $\nabla_{A}$ and $\nabla_{G}$ are initialized according to Eq. 8 and Eq. 10. When system events happen on the node, and tags are updated according to the propagation policies, the gradients are also updated. Theoretically, the worst case of memory usage for saving all $\nabla_{n}^{A}$ is $O(|N|^{2})$, and $O(|N||E|)$ for saving all $\nabla_{n}^{G}$.
@@ C-B Amount of Gradients to Save During Training
Please note that we only store non-zero gradients in the dictionary. Therefore, although $|N|$ and $|E|$ could be very large in provenance graphs, the size of $\nabla^{A}$ and $\nabla^{G}$ would be reasonable because gradients with the value of zero would not be stored. Every time after the gradients are calculated, we add the non-zero values into $\nabla_{n}^{A}$ and $\nabla_{n}^{G}$ and remove the too-small gradients (e.g., $<10^{-5}$) from the dictionaries. In practice, most gradients would always be zero due to the following three reasons: 1) In most cases, the provenance graph is relatively sparse; a majority of nodes lack connections to other nodes, and only a small fraction of nodes engage in numerous interactions with others. For example, a specific process would only read a small set of files on the system; 2) Even if a node connects with many different neighbors, the gradients only get updated when the edge (i.e., the corresponding event) changes the tags. Therefore, most benign nodes and edges would not be added to $\nabla_{n}^{A}$ and $\nabla_{n}^{G}$; 3) As $g_{e}$ is smaller than 1, some gradients would become near-zero after many iterations, which would be removed from $\nabla_{n}^{A}$ and $\nabla_{n}^{G}$.
We show the distribution of the number of the non-zero gradients that need to be stored during tag propagation in Fig. 8. We use datasets from DARPA Engagements in this experiment because the duration is much longer than others. The histograms clearly illustrate the long-tailed distribution of the non-zero values of gradients for each node, which means we only need to store a tiny number of non-zero gradients for most nodes. The average numbers are less than 5 for all those four datasets. Hence, the additional overhead is reasonable after adding the learning module.
@@
C-C Mimicry Attack Experiment
Since it’s hard to perform mimicry attack behaviors in the DARPA engagement environment, we insert events into the existing DARPA dataset to mimic benign behaviors involving malicious entities. we used the E3 cadets testing set used in FLASH evaluation for this experiment. In the attack, /tmp/test file is downloaded and executed, serving as a command and control malware to carry out the rest of the attack. We identified a benign file /dev/tty to serve as the target to mimic. Specifically, we replicated interactions with /dev/tty on /tmp/test to let /tmp/test mimic normal behaviors.
C-D Ablation Study on Learning Rate
Figure 9: Training loss using different learning rates (lr).
Fig. 9 shows the changes of the training loss when we use different learning rates during training. For this experiment, we used Cadets from DARPA Engagement 3. We can see the influence of the learning rates on the training process. In practice, if the training is limited to 20 epochs (due to time and resource issues), a learning rate of 0.01 is advisable; for longer training durations, 0.001 is likely a more preferred choice.

### v2

@@ VIII Conclusion
This paper introduces Captain, a rule-based PIDS capable of automatically adapting to detection environments, enhanced by three adaptive parameters: tag initialization ($A$), propagation rate ($G$), and threshold ($T$). The differentiable detection framework enables the optimization using the gradient descent algorithm. To our knowledge, this is the first effort to incorporate gradient descent methods in optimizing rule-based PIDS. We evaluated Captain on several datasets. The results demonstrate the superior detection capability, significantly reducing false alarms, detection latency, and runtime overhead, outperforming the SOTA baselines.
Acknowledgment
We would like to thank anonymous reviewers for their constructive feedback. Lingzhi and Xiangmin were supported by the National Science Foundation (NSF) grant 2148177 and funds from the Resilient & Intelligent NextG Systems (RINGS) program. In addition, Zhenyuan was supported by the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024C03288). Sekar’s work was supported in part by NSF grants 1918667 and 2153056.
@@ Appendix A Attack Scenarios in Datasets
For benign scenarios, the SOC collaborating with us provided three detailed operation environments in the real world: Streaming simulates a data store and streaming server; Dev simulates a development server; Cloud simulates a cloud server. The simulated APT attack chains are generated using the Atomic Red Team [54]. We designed three types of simulated attacks: fully randomized, partially randomized, and deterministic attacks. Each step in a fully randomized attack is randomly selected from the Atomic Red Team library. A partially randomized attack has determined steps at some stages but randomly chooses steps for the rest of the steps. A deterministic attack has defined steps at every attack stage. We generated one deterministic and three partially randomized attacks in our evaluation. We conduct the experiments following the attack campaign pattern in DARPA Engagements to make it as realistic as possible. Table VII specifies the apps or processes running in each simulated environment.
TABLE VII: Simulated Scenario Summary
Attack Attack Description
ReverseShell(RS) Connect to the victim’s host, collect system information, and install multiple applications
@@
Benign App or Process Involved
Streaming Kafka, Mysql, Nginx, Redis, Zookeeper
@@ Appendix B Captain Policies
We adopt tag propagation rules, detailed in Table VIII, from Morse [22] for unbiased comparison. $dtag$ refers to the data tag and $ptag$ refers to the code tag. Please refer to the original paper for formal and detailed descriptions of the hyperparameters such as $T_{qb}$, $T_{qe}$, $d_{b}$, and $d_{e}$. All hyperparameters are set according to the recommendation in [22]. We also adopt the alarm generation rules from Morse [22], detailed in Table IX. $incl\_exec(p)$ means $p$ is contains the execution permission. $socket(o)$ holds when $o$ refers to a socket.
TABLE VIII: Tag propagation policies
@@
benign suspect suspect environment
@@
TABLE IX: Alarm generation policies
@@
Appendix C Mathematical Proof of the Optimization
In this section, we prove the equations of gradients of $A$, $G$, and $T$ to the loss, i.e. Eq. 7, Eq. 9, and Eq. 10 in §IV-C. According to the loss function Eq.4 and the chain rule in calculus, we have:
$\frac{\partial Loss}{\partial a_{n}}$ $=\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}\cdot\frac{\partial f}{\partial a_{n}}+\alpha\cdot(a_{n}-a_{0})$ (13)
$\frac{\partial Loss}{\partial g_{e}}$ $=\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}\cdot\frac{\partial f}{\partial g_{e}}+\gamma\cdot(g_{e}-g_{0})$
$\frac{\partial Loss}{\partial thr_{e}}$ $=\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}\cdot\frac{\partial f}{\partial thr_{e}}+\tau\cdot(thr_{e}-thr_{0})$
where $\frac{\partial\mathcal{L}}{\partial f}$ is identical for all events according to Eq. 3 and $\alpha$, $\gamma$, and $\tau$ are pre-defined hyperparameters. We have
$f(e)$ $=tag_{dest}^{new}-thr_{e}$
$=g_{e}\cdot tag_{rule}+(1-g_{e})\cdot tag_{dest}-thr_{e}$
$\frac{\partial f}{\partial thr_{e}}$ $=-1$
$\frac{\partial f}{\partial a_{n}}$ $=g_{e}\cdot\frac{\partial tag_{rule}}{\partial a_{n}}+(1-g_{e})\cdot\frac{\partial tag_{dest}}{\partial a_{n}}$
$f(e)$ is influenced by the propagation rate of the current event (denoted by $g_{e}$) and all previous events (denoted by $g_{e^{\prime}}$). Consequently, every time when $tag_{dest}^{new}$ is updated, we recalculate $\frac{\partial f}{\partial g_{e}}$ and $\frac{\partial f}{\partial g_{e^{\prime}}}$.
$Since\ \frac{\partial g_{e}}{\partial g_{e^{\prime}}}$ $=0,$ (14)
$\frac{\partial f}{\partial g_{e^{\prime}}}$ $=g_{e}\cdot\frac{\partial tag_{rule}}{\partial g_{e^{\prime}}}+(1-g_{e})\cdot\frac{\partial tag_{dest}}{\partial g_{e^{\prime}}}$
The calculation of $\frac{\partial f}{\partial g_{e}}$ involves the product of two functions. According to the Product Rule:
$if\ h(x)=f(x)g(x),\ then\ h^{\prime}(x)=f^{\prime}(x)g(x)+f(x)g^{\prime}(x)$
we have
$\frac{\partial f}{\partial g_{e}}=g_{e}\frac{\partial tag_{rule}}{\partial g_{e}}+tag_{rule}+(1-g_{e})\frac{\partial tag_{dest}}{\partial g_{e}}-tag_{dest}$ (15)
Appendix D Mathematical Proof of the Heuristic Hyperparameter Setting
According to Eq.13, to let $\frac{\partial Loss}{\partial a_{n}}$, $\frac{\partial Loss}{\partial g_{e}}$, and $\frac{\partial Loss}{\partial thr_{e}}$ be zero, we have
$\alpha$ $=-\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}\cdot\frac{\partial f}{\partial a_{n}}\cdot(a_{n}-a_{0})^{-1}$ (16)
$\gamma$ $=-\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}\cdot\frac{\partial f}{\partial g_{e}}\cdot(g_{e}-g_{0})^{-1}$
$\tau$ $=-\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}\cdot\frac{\partial f}{\partial thr_{e}}\cdot(thr_{e}-thr_{0})^{-1}$
Since $\mathcal{L}(e)=max(0,(1-f(e))^{2}-1)$ and $-1<f(e)<0$,
$-\sum_{e\in E}\frac{\partial\mathcal{L}(e)}{\partial f}=\sum_{e\in E}2(1-f(e))$
$\frac{\partial f}{\partial thr_{e}}=-1$, $10<\frac{\partial f}{\partial a_{n}}<$ 1, $-1<\frac{\partial f}{\partial g_{e}}<0$. As mentioned in §IV-B, $a_{0}$ is usually 0 while $g_{0}$ is 1. We can make the following estimations:
$f(e)\approx-0.5,\partial f/\partial a_{n}\approx 0.5,\partial f/\partial g_{e}\approx-0.5$ (17)
$a_{n}-a_{0}\approx 0.5,g_{e}-g_{0}\approx-0.5,thr_{e}-thr_{0}\approx-0.25$
Let $N(e)$ be the number of the allowed false alarms related to event $e$ and $N(n)$ be the number of the allowed false alarms related to node $n$.
$\alpha$ $=\sum_{e\in E}2(1-f(e))\cdot\frac{\partial f}{\partial a_{n}}\cdot(a_{n}-a_{0})^{-1}\approx 3N(n)$ (18)
$\gamma$ $=\sum_{e\in E}2(1-f(e))\cdot\frac{\partial f}{\partial g_{e}}\cdot(g_{e}-g_{0})^{-1}\approx 3N(e)$
$\tau$ $=\sum_{e\in E}2(1-f(e))\cdot\frac{\partial f}{\partial thr_{e}}\cdot(thr_{e}-thr_{0})^{-1}\approx 12N(e)$
Appendix E Additional Experiment Results
E-A Gradients Storage During Training
@@
We maintain two dictionaries for each node $n$ to save the gradients of $tag_{n}$, $\nabla_{n}^{A}:\{n^{\prime}:\frac{\partial tag_{n}}{\partial a_{n}^{\prime}},n^{\prime}\in N\}$ and $\nabla_{n}^{G}:\{e:\frac{\partial tag_{n}}{\partial g_{e}},e\in E\}$. Theoretically, the worst case for saving $\nabla_{n}^{A}$ is $O(|N|^{2})$, and $O(|N||E|)$ for saving $\nabla_{n}^{G}$. Please note that we only store non-zero gradients. Every time after the gradients are calculated, we add the non-zero values into $\nabla_{n}^{A}$ and $\nabla_{n}^{G}$ and discard the too-small gradients (e.g., $<10^{-5}$). In practice, most gradients would always be zero for three reasons: 1) In most cases, the provenance graph is relatively sparse. For example, a specific process would only read a small set of files on the system; 2) The gradients only get updated when the event changes the tags. Therefore, most benign nodes and edges would not be added to $\nabla_{n}^{A}$ and $\nabla_{n}^{G}$; 3) As $g_{e}$ is smaller than 1, some gradients would become near-zero after many iterations. Fig. 8 clearly illustrates the long-tailed distribution of the non-zero gradients on the DARPA dataset, which means we only need to store a tiny number of non-zero gradients for most nodes. The average numbers are less than 5 for all those four datasets. Hence, the additional overhead is reasonable during training.
@@
E-B Mimicry Attack Experiment
We insert events into the dataset to mimic benign behaviors involving malicious entities. we used the E3 Cadets for this experiment. In the attack, /tmp/test file is downloaded and executed, serving as a command and control malware to carry out the rest of the attack. We identified a benign file /dev/tty to serve as the target to mimic and replicated interactions with /dev/tty on /tmp/test to let /tmp/test mimic normal behaviors.
E-C Ablation Study on Learning Rate
Figure 9: Training loss using different learning rates.
Fig. 9 shows the training loss with different learning rates on Cadets from DARPA Engagement 3. From the figure, if the training is limited to 20 epochs (due to time and resource issues), a learning rate of 0.01 is advisable; for longer training durations, 0.001 is likely a more preferred choice.